OpenStudy (anonymous):

a wooden block of mass 8 kg is tied to a string attatched to the bottom of the tank.In equilibrium the block is completely immersed in water. if relative density of wood is 0.8 and g=10m/s^2 then what is the value of tension in the string?

OpenStudy (anonymous):

any 1 pleeeese help

OpenStudy (anonymous):

hey jamesj pleese help

OpenStudy (jamesj):

There are three forces acting on the block of wood F_g = force of gravity (down) F_b = force of buoyancy (up) F_s = force from the string (down) The block is in equilibrium hence the sum of the three is zero. Use that to solve for F_s. The magnitude of F_s is the tension in the string.

OpenStudy (jamesj):

Make sense?

OpenStudy (anonymous):

wow i am extremely convinced

OpenStudy (anonymous):

but for more convinciation can u pleese solve the problem for me

OpenStudy (jamesj):

No ... I want you to take the next steps. What are expressions for \[ F_g \] and \[ F_b \] ?

OpenStudy (jamesj):

F_g is particularly easy

OpenStudy (anonymous):

is f_g is mg?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

so u dono the expressions?

OpenStudy (jamesj):

I do know. Of course. But I want you to learn by doing, not watching.

OpenStudy (anonymous):

k pleese wait

OpenStudy (jamesj):

To find F_b, go back to first principles and definitions The buoyant force is equal to the weight of water displaced. hence you need to find the weight of water displaced. The find the weight of water displaced, you'll need to find the volume of water displaced and multiply it by the density of water. Then to find the volume of water displaced ...

OpenStudy (anonymous):

buoyant force=weight of water displaced

OpenStudy (anonymous):

i.e F_b=mg

OpenStudy (jamesj):

Ok, yes, where m here is the mass of the water displaced

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so now can u solve

OpenStudy (jamesj):

We use rho, \( \rho \) for density. Hence you have by definition \[ \rho_{water} = \frac{m_{water}}{V_{water}} \] where V is volume. Hence \[ m_{water} = \rho_{water}.V_{water} \] Now find \( V_{water} \). It must be equal to \( V_{block} \). Hence you have to find \( V_{block} \). Use the same density principle. That's why the density of wood was given to you in the problem.

OpenStudy (anonymous):

so wat's the use of relative density

OpenStudy (jamesj):

Tell me first what's the equation of the volume of the block?

OpenStudy (anonymous):

of cousre it is m/d

OpenStudy (anonymous):

d is the density

OpenStudy (jamesj):

Hence \[ m_{water} = \rho_{water}.V_{water} \] \[ = \rho_{water} . V_{block} \] because the volume of water displaced must equal the volume of the block \[ = \rho_{water}.\frac{m_{block}}{\rho_{block}} \ \ \ \ \text{ by definition of density } \] \[ = \frac{\rho_{water}}{\rho_{block}}. m_{block} \] See what to do now?

OpenStudy (anonymous):

i cant understand the last equation

OpenStudy (anonymous):

k thanks for yor help iam really happy with the presence of u

OpenStudy (jamesj):

what don't you understand about the last equation.

OpenStudy (anonymous):

no other guy was able to answer my qusetion

OpenStudy (anonymous):

i'll send u a frnd request in face book

OpenStudy (anonymous):

what is your fb username and i am nt big enough to misuse it

OpenStudy (jamesj):

By definition \[ \rho_{block}/\rho_{water} \] is the relative density of the block, i.e., is equal to 0.8 Therefore \[ m_{water} = \rho_{water}/\rho_{block}.m_{block} \] \[ = (\rho_{block}/\rho_{water})^{-1}.m_block \] \[ = (0.8)^{-1}.m_block \] \[ = (5/4).m_block \] \[ = (5/4).(8 \ kg) \] \[ = 10 \ kg \] Now I want you to make an effort to finish the problem yourself, as I've now shown you 90% of it.

OpenStudy (jamesj):

I should have written \[ m_{block} \] not \[ m_block \]

OpenStudy (anonymous):

wow wow wo o

OpenStudy (jamesj):

I'm going to help someone else with a question. Try and figure it out while I'm gone. I'll keep an eye on what's going on over here.

OpenStudy (anonymous):

k by the way pleese tell yor fb username

OpenStudy (jamesj):

No, I don't do that, sorry.